How many grams of glucose are needed to prepare 552 mL of a 2.0%(m/v) glucose solution?
Answer #1

Mass/Volume% = (mass of solute/volume of solvent)*100

mass of solute = (Mass/Volume%)*(volue of solvent)/100

= (2*552)/100 = 11.04 grams

mass of solute needed to make 2%(m/v) glucose solution = 11.04 grams